### v1

Jilei Xu
xujl@ihep.ac.cn
Institute of High Energy Physics, Beijing, China
Miao He
hem@ihep.ac.cn
Institute of High Energy Physics, Beijing, China
Cédric Cerna
cedric.cerna@in2p3.fr
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Yongbo Huang
huangyb@gxu.edu.cn
Guangxi University, Nanning, China
Thomas Adam
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Shakeel Ahmad
Pakistan Institute of Nuclear Science and Technology, Islamabad, Pakistan
Rizwan Ahmed
Pakistan Institute of Nuclear Science and Technology, Islamabad, Pakistan
Fengpeng An
Sun Yat-Sen University, Guangzhou, China
Costas Andreopoulos
The University of Liverpool, Department of Physics, Oliver Lodge Laboratory, Oxford Str., Liverpool L69 7ZE, UK, United Kingdom
Giuseppe Andronico
INFN Catania and Dipartimento di Fisica e Astronomia dell Università di Catania, Catania, Italy
João Pedro Athayde Marcondes de André
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Nikolay Anfimov
Joint Institute for Nuclear Research, Dubna, Russia
Vito Antonelli
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Tatiana Antoshkina
Joint Institute for Nuclear Research, Dubna, Russia
Didier Auguste
IJCLab, Université Paris-Saclay, CNRS/IN2P3, 91405 Orsay, France
Weidong Bai
Sun Yat-Sen University, Guangzhou, China
Nikita Balashov
Joint Institute for Nuclear Research, Dubna, Russia
Andrea Barresi
INFN Milano Bicocca and University of Milano Bicocca, Milano, Italy
Davide Basilico
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Eric Baussan
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Marco Beretta
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Antonio Bergnoli
INFN Sezione di Padova, Padova, Italy
Nikita Bessonov
Joint Institute for Nuclear Research, Dubna, Russia
Daniel Bick
Institute of Experimental Physics, University of Hamburg, Hamburg, Germany
Lukas Bieger
Eberhard Karls Universität Tübingen, Physikalisches Institut, Tübingen, Germany
Svetlana Biktemerova
Joint Institute for Nuclear Research, Dubna, Russia
Thilo Birkenfeld
III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Simon Blyth
Institute of High Energy Physics, Beijing, China
Anastasia Bolshakova
Joint Institute for Nuclear Research, Dubna, Russia
Mathieu Bongrand
Nantes Université, IMT Atlantique, CNRS/IN2P3, Nantes, France
Matteo Borghesi
INFN Milano Bicocca and University of Milano Bicocca, Milano, Italy
Dominique Breton
IJCLab, Université Paris-Saclay, CNRS/IN2P3, 91405 Orsay, France
Augusto Brigatti
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Riccardo Brugnera
Dipartimento di Fisica e Astronomia dell’Università di Padova and INFN Sezione di Padova, Padova, Italy
Riccardo Bruno
INFN Catania and Dipartimento di Fisica e Astronomia dell Università di Catania, Catania, Italy
Antonio Budano
Dipartimento di Matematica e Fisica, Università Roma Tre and INFN Sezione Roma Tre, Roma, Italy
Jose Busto
Aix Marseille Univ, CNRS/IN2P3, CPPM, Marseille, France
Marcel Büchner
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Anatael Cabrera
IJCLab, Université Paris-Saclay, CNRS/IN2P3, 91405 Orsay, France
Barbara Caccianiga
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Hao Cai
Wuhan University, Wuhan, China
Xiao Cai
Institute of High Energy Physics, Beijing, China
Yanke Cai
Institute of High Energy Physics, Beijing, China
Zhiyan Cai
Institute of High Energy Physics, Beijing, China
Stéphane Callier
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Steven Calvez
Nantes Université, IMT Atlantique, CNRS/IN2P3, Nantes, France
Antonio Cammi
Now at Emirates Nuclear Technology Center (ENTC), Khalifa University, Abu Dhabi, United Arab Emirates.
INFN Milano Bicocca and Politecnico of Milano, Milano, Italy
Chuanya Cao
Institute of High Energy Physics, Beijing, China
Guofu Cao
Institute of High Energy Physics, Beijing, China
Jun Cao
Institute of High Energy Physics, Beijing, China
Yaoqi Cao
University of Warwick, University of Warwick, Coventry, CV4 7AL, United Kingdom
The University of Liverpool, Department of Physics, Oliver Lodge Laboratory, Oxford Str., Liverpool L69 7ZE, UK, United Kingdom
Rossella Caruso
INFN Catania and Dipartimento di Fisica e Astronomia dell Università di Catania, Catania, Italy
Vanessa Cerrone
Dipartimento di Fisica e Astronomia dell’Università di Padova and INFN Sezione di Padova, Padova, Italy
Jinfan Chang
Institute of High Energy Physics, Beijing, China
Yun Chang
National United University, Miao-Li
Auttakit Chatrabhuti
High Energy Physics Research Unit, Faculty of Science, Chulalongkorn University, Bangkok, Thailand
Chao Chen
Institute of High Energy Physics, Beijing, China
Guoming Chen
Guangxi University, Nanning, China
Jiahui Chen
Institute of High Energy Physics, Beijing, China
Jian Chen
Sun Yat-Sen University, Guangzhou, China
Jing Chen
Sun Yat-Sen University, Guangzhou, China
Junyou Chen
Guangxi University, Nanning, China
Pingping Chen
Dongguan University of Technology, Dongguan, China
Shaomin Chen
Tsinghua University, Beijing, China
Shiqiang Chen
Nanjing University, Nanjing, China
Xin Chen
Nanjing University, Nanjing, China
Institute of High Energy Physics, Beijing, China
Yiming Chen
Institute of High Energy Physics, Beijing, China
Yixue Chen
North China Electric Power University, Beijing, China
Yu Chen
Sun Yat-Sen University, Guangzhou, China
Ze Chen
GSI Helmholtzzentrum für Schwerionenforschung GmbH, Planckstr. 1, 64291 Darmstadt (Germany)
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Zhangming Chen
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Zhiyuan Chen
Institute of High Energy Physics, Beijing, China
Jie Cheng
North China Electric Power University, Beijing, China
Yaping Cheng
Beijing Institute of Spacecraft Environment Engineering, Beijing, China
Yu Chin Cheng
Department of Physics, National Taiwan University, Taipei
Alexander Chepurnov
Lomonosov Moscow State University, Moscow, Russia
Alexey Chetverikov
Joint Institute for Nuclear Research, Dubna, Russia
Davide Chiesa
INFN Milano Bicocca and University of Milano Bicocca, Milano, Italy
Pietro Chimenti
Universidade Estadual de Londrina, Londrina, Brazil
Po-Lin Chou
Institute of Physics, National Yang Ming Chiao Tung University, Hsinchu
Ziliang Chu
Institute of High Energy Physics, Beijing, China
Artem Chukanov
Joint Institute for Nuclear Research, Dubna, Russia
Gérard Claverie
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Catia Clementi
INFN Sezione di Perugia and Dipartimento di Chimica, Biologia e Biotecnologie dell’Università di Perugia, Perugia, Italy
Barbara Clerbaux
Université Libre de Bruxelles, Brussels, Belgium
Claudio Coletta
INFN Milano Bicocca and University of Milano Bicocca, Milano, Italy
Simon Csakli
Technische Universität München, München, Germany
Chenyang Cui
Institute of High Energy Physics, Beijing, China
Olivia Dalager
Department of Physics and Astronomy, University of California, Irvine, California, USA
Zhi Deng
Tsinghua University, Beijing, China
Ziyan Deng
Institute of High Energy Physics, Beijing, China
Xiaoyu Ding
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Xuefeng Ding
Institute of High Energy Physics, Beijing, China
Yayun Ding
Institute of High Energy Physics, Beijing, China
Bayu Dirgantara
Suranaree University of Technology, Nakhon Ratchasima, Thailand
Carsten Dittrich
Technische Universität München, München, Germany
Sergey Dmitrievsky
Joint Institute for Nuclear Research, Dubna, Russia
David Doerflinger
Technische Universität München, München, Germany
Dmitry Dolzhikov
Joint Institute for Nuclear Research, Dubna, Russia
Haojie Dong
Institute of High Energy Physics, Beijing, China
Jianmeng Dong
Tsinghua University, Beijing, China
Evgeny Doroshkevich
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Marcos Dracos
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Frédéric Druillole
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Ran Du
Institute of High Energy Physics, Beijing, China
Shuxian Du
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Yujie Duan
Wuhan University, Wuhan, China
Katherine Dugas
Department of Physics and Astronomy, University of California, Irvine, California, USA
Stefano Dusini
INFN Sezione di Padova, Padova, Italy
Hongyue Duyang
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Jessica Eck
Eberhard Karls Universität Tübingen, Physikalisches Institut, Tübingen, Germany
Timo Enqvist
University of Jyvaskyla, Department of Physics, Jyvaskyla, Finland
Andrea Fabbri
Dipartimento di Matematica e Fisica, Università Roma Tre and INFN Sezione Roma Tre, Roma, Italy
Ulrike Fahrendholz
Technische Universität München, München, Germany
Lei Fan
Institute of High Energy Physics, Beijing, China
Jian Fang
Institute of High Energy Physics, Beijing, China
Wenxing Fang
Institute of High Energy Physics, Beijing, China
Elia Stanescu Farilla
Dipartimento di Matematica e Fisica, Università Roma Tre and INFN Sezione Roma Tre, Roma, Italy
Dmitry Fedoseev
Joint Institute for Nuclear Research, Dubna, Russia
Li-Cheng Feng
Institute of Physics, National Yang Ming Chiao Tung University, Hsinchu
Qichun Feng
Harbin Institute of Technology, Harbin, China
Federico Ferraro
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Daniela Fetzer
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Marcellin Fotzé
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Amélie Fournier
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Aaron Freegard
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Feng Gao
Université Libre de Bruxelles, Brussels, Belgium
Alberto Garfagnini
Dipartimento di Fisica e Astronomia dell’Università di Padova and INFN Sezione di Padova, Padova, Italy
Arsenii Gavrikov
Dipartimento di Fisica e Astronomia dell’Università di Padova and INFN Sezione di Padova, Padova, Italy
Marco Giammarchi
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Nunzio Giudice
INFN Catania and Dipartimento di Fisica e Astronomia dell Università di Catania, Catania, Italy
Maxim Gonchar
Joint Institute for Nuclear Research, Dubna, Russia
Guanghua Gong
Tsinghua University, Beijing, China
Hui Gong
Tsinghua University, Beijing, China
Yuri Gornushkin
Joint Institute for Nuclear Research, Dubna, Russia
Marco Grassi
Dipartimento di Fisica e Astronomia dell’Università di Padova and INFN Sezione di Padova, Padova, Italy
Maxim Gromov
Lomonosov Moscow State University, Moscow, Russia
Vasily Gromov
Joint Institute for Nuclear Research, Dubna, Russia
Minhao Gu
Institute of High Energy Physics, Beijing, China
Xiaofei Gu
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Yu Gu
Jinan University, Guangzhou, China
Mengyun Guan
Institute of High Energy Physics, Beijing, China
Yuduo Guan
Institute of High Energy Physics, Beijing, China
Nunzio Guardone
INFN Catania and Dipartimento di Fisica e Astronomia dell Università di Catania, Catania, Italy
Rosa Maria Guizzetti
Dipartimento di Fisica e Astronomia dell’Università di Padova and INFN Sezione di Padova, Padova, Italy
Cong Guo
Institute of High Energy Physics, Beijing, China
Wanlei Guo
Institute of High Energy Physics, Beijing, China
Caren Hagner
Institute of Experimental Physics, University of Hamburg, Hamburg, Germany
Hechong Han
Institute of High Energy Physics, Beijing, China
Ran Han
Beijing Institute of Spacecraft Environment Engineering, Beijing, China
Yang Han
Sun Yat-Sen University, Guangzhou, China
Vidhya Thara Hariharan
Institute of Experimental Physics, University of Hamburg, Hamburg, Germany
Jinhong He
Wuhan University, Wuhan, China
Wei He
Institute of High Energy Physics, Beijing, China
Xinhai He
Institute of High Energy Physics, Beijing, China
Ziou He
University of Warwick, University of Warwick, Coventry, CV4 7AL, United Kingdom
The University of Liverpool, Department of Physics, Oliver Lodge Laboratory, Oxford Str., Liverpool L69 7ZE, UK, United Kingdom
Tobias Heinz
Eberhard Karls Universität Tübingen, Physikalisches Institut, Tübingen, Germany
Patrick Hellmuth
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Yuekun Heng
Institute of High Energy Physics, Beijing, China
YuenKeung Hor
Sun Yat-Sen University, Guangzhou, China
Shaojing Hou
Institute of High Energy Physics, Beijing, China
Yee Hsiung
Department of Physics, National Taiwan University, Taipei
Bei-Zhen Hu
Department of Electro-Optical Engineering, National Taipei University of Technology, Taipei
Hang Hu
Sun Yat-Sen University, Guangzhou, China
Jun Hu
Institute of High Energy Physics, Beijing, China
Tao Hu
Institute of High Energy Physics, Beijing, China
Yuxiang Hu
Institute of High Energy Physics, Beijing, China
Guihong Huang
Wuyi University, Jiangmen, China
Hexiang Huang
Institute of High Energy Physics, Beijing, China
Jinhao Huang
Institute of High Energy Physics, Beijing, China
Junting Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Kaixuan Huang
Sun Yat-Sen University, Guangzhou, China
Shengheng Huang
Wuyi University, Jiangmen, China
Tao Huang
Sun Yat-Sen University, Guangzhou, China
Xin Huang
Institute of High Energy Physics, Beijing, China
Xingtao Huang
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Jiaqi Hui
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Lei Huo
Harbin Institute of Technology, Harbin, China
Cédric Huss
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Safeer Hussain
Pakistan Institute of Nuclear Science and Technology, Islamabad, Pakistan
Leonard Imbert
Nantes Université, IMT Atlantique, CNRS/IN2P3, Nantes, France
Ara Ioannisian
Yerevan Physics Institute, Yerevan, Armenia
Adrienne Jacobi
Department of Physics and Astronomy, University of California, Irvine, California, USA
Arshak Jafar
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Beatrice Jelmini
Dipartimento di Fisica e Astronomia dell’Università di Padova and INFN Sezione di Padova, Padova, Italy
Xiangpan Ji
Nankai University, Tianjin, China
Xiaolu Ji
Institute of High Energy Physics, Beijing, China
Huihui Jia
Nankai University, Tianjin, China
Junji Jia
Wuhan University, Wuhan, China
Cailian Jiang
Nanjing University, Nanjing, China
Wei Jiang
Institute of High Energy Physics, Beijing, China
Xiaoshan Jiang
Institute of High Energy Physics, Beijing, China
Xiaozhao Jiang
Institute of High Energy Physics, Beijing, China
Yijian Jiang
North China Electric Power University, Beijing, China
Yixuan Jiang
Institute of High Energy Physics, Beijing, China
Xiaoping Jing
Institute of High Energy Physics, Beijing, China
Cécile Jollet
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Li Kang
Dongguan University of Technology, Dongguan, China
Rebin Karaparabil
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Narine Kazarian
Yerevan Physics Institute, Yerevan, Armenia
Ali Khan
Pakistan Institute of Nuclear Science and Technology, Islamabad, Pakistan
Amina Khatun
Université Libre de Bruxelles, Brussels, Belgium
Comenius University Bratislava, Faculty of Mathematics, Physics and Informatics, Bratislava, Slovakia
Khanchai Khosonthongkee
Suranaree University of Technology, Nakhon Ratchasima, Thailand
Denis Korablev
Joint Institute for Nuclear Research, Dubna, Russia
Konstantin Kouzakov
Lomonosov Moscow State University, Moscow, Russia
Alexey Krasnoperov
Joint Institute for Nuclear Research, Dubna, Russia
Sergey Kuleshov
Millennium Institute for SubAtomic Physics at the High-energy Frontier (SAPHIR), ANID, Chile
Sindhujha Kumaran
Department of Physics and Astronomy, University of California, Irvine, California, USA
Nikolay Kutovskiy
Joint Institute for Nuclear Research, Dubna, Russia
Loïc Labit
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Tobias Lachenmaier
Eberhard Karls Universität Tübingen, Physikalisches Institut, Tübingen, Germany
Haojing Lai
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Cecilia Landini
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Lorenzo Lastrucci
INFN Sezione di Padova, Padova, Italy
Sébastien Leblanc
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Matthieu Lecocq
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Frederic Lefevre
Nantes Université, IMT Atlantique, CNRS/IN2P3, Nantes, France
Ruiting Lei
Dongguan University of Technology, Dongguan, China
Rupert Leitner
Charles University, Faculty of Mathematics and Physics, Prague, Czech Republic
Jason Leung
Institute of Physics, National Yang Ming Chiao Tung University, Hsinchu
Demin Li
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Fei Li
Institute of High Energy Physics, Beijing, China
Fule Li
Tsinghua University, Beijing, China
Gaosong Li
Institute of High Energy Physics, Beijing, China
Hongjian Li
Institute of High Energy Physics, Beijing, China
Huang Li
Institute of High Energy Physics, Beijing, China
Jiajun Li
Sun Yat-Sen University, Guangzhou, China
Min Li
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Nan Li
College of Electronic Science and Engineering, National University of Defense Technology, Changsha, China
Qingjiang Li
College of Electronic Science and Engineering, National University of Defense Technology, Changsha, China
Ruhui Li
Institute of High Energy Physics, Beijing, China
Rui Li
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Shanfeng Li
Dongguan University of Technology, Dongguan, China
Tao Li
Sun Yat-Sen University, Guangzhou, China
Teng Li
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Weidong Li
Institute of High Energy Physics, Beijing, China
University of Chinese Academy of Sciences, Beijing, China
Xiaonan Li
Institute of High Energy Physics, Beijing, China
Yi Li
Dongguan University of Technology, Dongguan, China
Yichen Li
Institute of High Energy Physics, Beijing, China
Yifan Li
Institute of High Energy Physics, Beijing, China
Yufeng Li
Institute of High Energy Physics, Beijing, China
Zhaohan Li
Institute of High Energy Physics, Beijing, China
Zhibing Li
Sun Yat-Sen University, Guangzhou, China
Zi-Ming Li
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Zonghai Li
Wuhan University, Wuhan, China
An-An Liang
Institute of Physics, National Yang Ming Chiao Tung University, Hsinchu
Jiajun Liao
Sun Yat-Sen University, Guangzhou, China
Minghua Liao
Sun Yat-Sen University, Guangzhou, China
Yilin Liao
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Ayut Limphirat
Suranaree University of Technology, Nakhon Ratchasima, Thailand
Bo-Chun Lin
Institute of Physics, National Yang Ming Chiao Tung University, Hsinchu
Guey-Lin Lin
Institute of Physics, National Yang Ming Chiao Tung University, Hsinchu
Shengxin Lin
Dongguan University of Technology, Dongguan, China
Tao Lin
Institute of High Energy Physics, Beijing, China
Xianhao Lin
Nanjing University, Nanjing, China
Xingyi Lin
Guangxi University, Nanning, China
Jiajie Ling
Sun Yat-Sen University, Guangzhou, China
Xin Ling
University of South China, Hengyang, China
Ivano Lippi
INFN Sezione di Padova, Padova, Italy
Caimei Liu
Institute of High Energy Physics, Beijing, China
Fang Liu
North China Electric Power University, Beijing, China
Fengcheng Liu
North China Electric Power University, Beijing, China
Haidong Liu
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Haotian Liu
Wuhan University, Wuhan, China
Hongbang Liu
Guangxi University, Nanning, China
Hongjuan Liu
University of South China, Hengyang, China
Hongtao Liu
Sun Yat-Sen University, Guangzhou, China
Hongyang Liu
Institute of High Energy Physics, Beijing, China
Jianglai Liu
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, China
Jiaxi Liu
Institute of High Energy Physics, Beijing, China
Jinchang Liu
Institute of High Energy Physics, Beijing, China
Kainan Liu
Wuyi University, Jiangmen, China
Min Liu
University of South China, Hengyang, China
Qian Liu
University of Chinese Academy of Sciences, Beijing, China
Runxuan Liu
GSI Helmholtzzentrum für Schwerionenforschung GmbH, Planckstr. 1, 64291 Darmstadt (Germany)
III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Shenghui Liu
Institute of High Energy Physics, Beijing, China
Shulin Liu
Institute of High Energy Physics, Beijing, China
Xiaowei Liu
Sun Yat-Sen University, Guangzhou, China
Xiwen Liu
Guangxi University, Nanning, China
Xuewei Liu
Tsinghua University, Beijing, China
Yankai Liu
Xi’an Jiaotong University, Xi’an, China
Zhen Liu
Institute of High Energy Physics, Beijing, China
Lorenzo Loi
INFN Milano Bicocca and University of Milano Bicocca, Milano, Italy
Alexey Lokhov
Lomonosov Moscow State University, Moscow, Russia
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Paolo Lombardi
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Claudio Lombardo
INFN Catania and Dipartimento di Fisica e Astronomia dell Università di Catania, Catania, Italy
Kai Loo
University of Jyvaskyla, Department of Physics, Jyvaskyla, Finland
Selma Conforti Di Lorenzo
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Haoqi Lu
Institute of High Energy Physics, Beijing, China
Junguang Lu
Institute of High Energy Physics, Beijing, China
Meishu Lu
Technische Universität München, München, Germany
Shuxiang Lu
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Xianguo Lu
University of Warwick, University of Warwick, Coventry, CV4 7AL, United Kingdom
Bayarto Lubsandorzhiev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Sultim Lubsandorzhiev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Livia Ludhova
GSI Helmholtzzentrum für Schwerionenforschung GmbH, Planckstr. 1, 64291 Darmstadt (Germany)
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Arslan Lukanov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Fengjiao Luo
University of South China, Hengyang, China
Guang Luo
Sun Yat-Sen University, Guangzhou, China
Jianyi Luo
Sun Yat-Sen University, Guangzhou, China
Shu Luo
Xiamen University, Xiamen, China
Wuming Luo
Institute of High Energy Physics, Beijing, China
Xiaojie Luo
Institute of High Energy Physics, Beijing, China
Vladimir Lyashuk
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Bangzheng Ma
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Bing Ma
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Qiumei Ma
Institute of High Energy Physics, Beijing, China
Si Ma
Institute of High Energy Physics, Beijing, China
Wing Yan Ma
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Xiaoyan Ma
Institute of High Energy Physics, Beijing, China
Xubo Ma
North China Electric Power University, Beijing, China
Jihane Maalmi
IJCLab, Université Paris-Saclay, CNRS/IN2P3, 91405 Orsay, France
Jingyu Mai
Sun Yat-Sen University, Guangzhou, China
Marco Malabarba
GSI Helmholtzzentrum für Schwerionenforschung GmbH, Planckstr. 1, 64291 Darmstadt (Germany)
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Yury Malyshkin
GSI Helmholtzzentrum für Schwerionenforschung GmbH, Planckstr. 1, 64291 Darmstadt (Germany)
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Roberto Carlos Mandujano
Department of Physics and Astronomy, University of California, Irvine, California, USA
Fabio Mantovani
Department of Physics and Earth Science, University of Ferrara and INFN Sezione di Ferrara, Ferrara, Italy
Xin Mao
Beijing Institute of Spacecraft Environment Engineering, Beijing, China
Stefano M. Mari
Dipartimento di Matematica e Fisica, Università Roma Tre and INFN Sezione Roma Tre, Roma, Italy
Agnese Martini
Laboratori Nazionali di Frascati dell’INFN, Roma, Italy
Matthias Mayer
Technische Universität München, München, Germany
Davit Mayilyan
Yerevan Physics Institute, Yerevan, Armenia
Yue Meng
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Anselmo Meregaglia
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Lino Miramonti
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Marta Colomer Molla
Université Libre de Bruxelles, Brussels, Belgium
Michele Montuschi
Department of Physics and Earth Science, University of Ferrara and INFN Sezione di Ferrara, Ferrara, Italy
Iwan Morton-Blake
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, China
Massimiliano Nastasi
INFN Milano Bicocca and University of Milano Bicocca, Milano, Italy
Dmitry V. Naumov
Joint Institute for Nuclear Research, Dubna, Russia
Elena Naumova
Joint Institute for Nuclear Research, Dubna, Russia
Igor Nemchenok
Joint Institute for Nuclear Research, Dubna, Russia
Elisabeth Neuerburg
III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Alexey Nikolaev
Lomonosov Moscow State University, Moscow, Russia
Feipeng Ning
Institute of High Energy Physics, Beijing, China
Zhe Ning
Institute of High Energy Physics, Beijing, China
Yujie Niu
Institute of High Energy Physics, Beijing, China
Hiroshi Nunokawa
Pontificia Universidade Catolica do Rio de Janeiro, Rio de Janeiro, Brazil
Lothar Oberauer
Technische Universität München, München, Germany
Juan Pedro Ochoa-Ricoux
Department of Physics and Astronomy, University of California, Irvine, California, USA
Millennium Institute for SubAtomic Physics at the High-energy Frontier (SAPHIR), ANID, Chile
Sebastian Olivares
Universidad Andres Bello, Fernandez Concha 700, Chile
Alexander Olshevskiy
Joint Institute for Nuclear Research, Dubna, Russia
Domizia Orestano
Dipartimento di Matematica e Fisica, Università Roma Tre and INFN Sezione Roma Tre, Roma, Italy
Fausto Ortica
INFN Sezione di Perugia and Dipartimento di Chimica, Biologia e Biotecnologie dell’Università di Perugia, Perugia, Italy
Rainer Othegraven
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Yifei Pan
Sun Yat-Sen University, Guangzhou, China
Alessandro Paoloni
Laboratori Nazionali di Frascati dell’INFN, Roma, Italy
George Parker
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Yatian Pei
Institute of High Energy Physics, Beijing, China
Luca Pelicci
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Anguo Peng
University of South China, Hengyang, China
Yu Peng
Institute of High Energy Physics, Beijing, China
Zhaoyuan Peng
Institute of High Energy Physics, Beijing, China
Elisa Percalli
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Willy Perrin
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Frédéric Perrot
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Pierre-Alexandre Petitjean
Université Libre de Bruxelles, Brussels, Belgium
Fabrizio Petrucci
Dipartimento di Matematica e Fisica, Università Roma Tre and INFN Sezione Roma Tre, Roma, Italy
Oliver Pilarczyk
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Artyom Popov
Lomonosov Moscow State University, Moscow, Russia
Pascal Poussot
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Ezio Previtali
INFN Milano Bicocca and University of Milano Bicocca, Milano, Italy
Fazhi Qi
Institute of High Energy Physics, Beijing, China
Ming Qi
Nanjing University, Nanjing, China
Xiaohui Qi
Institute of High Energy Physics, Beijing, China
Sen Qian
Institute of High Energy Physics, Beijing, China
Xiaohui Qian
Institute of High Energy Physics, Beijing, China
Zhonghua Qin
Institute of High Energy Physics, Beijing, China
Shoukang Qiu
University of South China, Hengyang, China
Manhao Qu
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Zhenning Qu
Institute of High Energy Physics, Beijing, China
Gioacchino Ranucci
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Thomas Raymond
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Alessandra Re
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Abdel Rebii
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Mariia Redchuk
INFN Sezione di Padova, Padova, Italy
Bin Ren
Dongguan University of Technology, Dongguan, China
Yuhan Ren
Institute of High Energy Physics, Beijing, China
Cristobal Morales Reveco
GSI Helmholtzzentrum für Schwerionenforschung GmbH, Planckstr. 1, 64291 Darmstadt (Germany)
III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Barbara Ricci
Department of Physics and Earth Science, University of Ferrara and INFN Sezione di Ferrara, Ferrara, Italy
Komkrit Rientong
High Energy Physics Research Unit, Faculty of Science, Chulalongkorn University, Bangkok, Thailand
Mariam Rifai
GSI Helmholtzzentrum für Schwerionenforschung GmbH, Planckstr. 1, 64291 Darmstadt (Germany)
III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Mathieu Roche
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Narongkiat Rodphai
Institute of High Energy Physics, Beijing, China
Fernanda de Faria Rodrigues
Institute of High Energy Physics, Beijing, China
Aldo Romani
INFN Sezione di Perugia and Dipartimento di Chimica, Biologia e Biotecnologie dell’Università di Perugia, Perugia, Italy
Bedřich Roskovec
Charles University, Faculty of Mathematics and Physics, Prague, Czech Republic
Arseniy Rybnikov
Joint Institute for Nuclear Research, Dubna, Russia
Andrey Sadovsky
Joint Institute for Nuclear Research, Dubna, Russia
Paolo Saggese
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Deshan Sandanayake
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Anut Sangka
National Astronomical Research Institute of Thailand, Chiang Mai, Thailand
Ujwal Santhosh
GSI Helmholtzzentrum für Schwerionenforschung GmbH, Planckstr. 1, 64291 Darmstadt (Germany)
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Giuseppe Sava
INFN Catania and Dipartimento di Fisica e Astronomia dell Università di Catania, Catania, Italy
Utane Sawangwit
National Astronomical Research Institute of Thailand, Chiang Mai, Thailand
Michaela Schever
III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Cédric Schwab
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Konstantin Schweizer
Technische Universität München, München, Germany
Alexandr Selyunin
Joint Institute for Nuclear Research, Dubna, Russia
Andrea Serafini
Dipartimento di Fisica e Astronomia dell’Università di Padova and INFN Sezione di Padova, Padova, Italy
Mariangela Settimo
Nantes Université, IMT Atlantique, CNRS/IN2P3, Nantes, France
Junyu Shao
Institute of High Energy Physics, Beijing, China
Vladislav Sharov
Joint Institute for Nuclear Research, Dubna, Russia
Hangyu Shi
Sun Yat-Sen University, Guangzhou, China
Hexi Shi
Dipartimento di Matematica e Fisica, Università Roma Tre and INFN Sezione Roma Tre, Roma, Italy
Jingyan Shi
Institute of High Energy Physics, Beijing, China
Yanan Shi
Institute of High Energy Physics, Beijing, China
Vitaly Shutov
Joint Institute for Nuclear Research, Dubna, Russia
Andrey Sidorenkov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Apeksha Singhal
GSI Helmholtzzentrum für Schwerionenforschung GmbH, Planckstr. 1, 64291 Darmstadt (Germany)
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Chiara Sirignano
Dipartimento di Fisica e Astronomia dell’Università di Padova and INFN Sezione di Padova, Padova, Italy
Jaruchit Siripak
Suranaree University of Technology, Nakhon Ratchasima, Thailand
Monica Sisti
INFN Milano Bicocca and University of Milano Bicocca, Milano, Italy
Mikhail Smirnov
Institute of Experimental Physics, University of Hamburg, Hamburg, Germany
Oleg Smirnov
Joint Institute for Nuclear Research, Dubna, Russia
Thiago Sogo-Bezerra
Nantes Université, IMT Atlantique, CNRS/IN2P3, Nantes, France
Sergey Sokolov
Joint Institute for Nuclear Research, Dubna, Russia
Julanan Songwadhana
Suranaree University of Technology, Nakhon Ratchasima, Thailand
Boonrucksar Soonthornthum
National Astronomical Research Institute of Thailand, Chiang Mai, Thailand
Albert Sotnikov
Joint Institute for Nuclear Research, Dubna, Russia
Warintorn Sreethawong
Suranaree University of Technology, Nakhon Ratchasima, Thailand
Achim Stahl
III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Luca Stanco
INFN Sezione di Padova, Padova, Italy
Konstantin Stankevich
Lomonosov Moscow State University, Moscow, Russia
Hans Steiger
Technische Universität München, München, Germany
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Jochen Steinmann
III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Tobias Sterr
Eberhard Karls Universität Tübingen, Physikalisches Institut, Tübingen, Germany
Matthias Raphael Stock
Technische Universität München, München, Germany
Virginia Strati
Department of Physics and Earth Science, University of Ferrara and INFN Sezione di Ferrara, Ferrara, Italy
Mikhail Strizh
Lomonosov Moscow State University, Moscow, Russia
Alexander Studenikin
Lomonosov Moscow State University, Moscow, Russia
Aoqi Su
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Jun Su
Sun Yat-Sen University, Guangzhou, China
Guangbao Sun
Wuhan University, Wuhan, China
Mingxia Sun
Institute of High Energy Physics, Beijing, China
Xilei Sun
Institute of High Energy Physics, Beijing, China
Yongzhao Sun
Institute of High Energy Physics, Beijing, China
Zhengyang Sun
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, China
Narumon Suwonjandee
High Energy Physics Research Unit, Faculty of Science, Chulalongkorn University, Bangkok, Thailand
Fedor Šimkovic
Comenius University Bratislava, Faculty of Mathematics, Physics and Informatics, Bratislava, Slovakia
Christophe De La Taille
Univ. Bordeaux, CNRS, LP2I, UMR 5797, F-33170 Gradignan, F-33170 Gradignan, France
Akira Takenaka
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, China
Xiaohan Tan
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Jian Tang
Sun Yat-Sen University, Guangzhou, China
Jingzhe Tang
Guangxi University, Nanning, China
Qiang Tang
Sun Yat-Sen University, Guangzhou, China
Quan Tang
University of South China, Hengyang, China
Xiao Tang
Institute of High Energy Physics, Beijing, China
Minh Thuan Nguyen Thi
Institute of Physics, National Yang Ming Chiao Tung University, Hsinchu
Yuxin Tian
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, China
Igor Tkachev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Tomas Tmej
Charles University, Faculty of Mathematics and Physics, Prague, Czech Republic
Marco Danilo Claudio Torri
INFN Sezione di Milano and Dipartimento di Fisica dell Università di Milano, Milano, Italy
Andrea Triossi
Dipartimento di Fisica e Astronomia dell’Università di Padova and INFN Sezione di Padova, Padova, Italy
Wladyslaw Trzaska
University of Jyvaskyla, Department of Physics, Jyvaskyla, Finland
Yu-Chen Tung
Department of Physics, National Kaohsiung Normal University, Kaohsiung
Cristina Tuve
INFN Catania and Dipartimento di Fisica e Astronomia dell Università di Catania, Catania, Italy
Nikita Ushakov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Carlo Venettacci
Dipartimento di Matematica e Fisica, Università Roma Tre and INFN Sezione Roma Tre, Roma, Italy
Giuseppe Verde
INFN Catania and Dipartimento di Fisica e Astronomia dell Università di Catania, Catania, Italy
Maxim Vialkov
Lomonosov Moscow State University, Moscow, Russia
Benoit Viaud
Nantes Université, IMT Atlantique, CNRS/IN2P3, Nantes, France
Cornelius Moritz Vollbrecht
GSI Helmholtzzentrum für Schwerionenforschung GmbH, Planckstr. 1, 64291 Darmstadt (Germany)
III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Vit Vorobel
Charles University, Faculty of Mathematics and Physics, Prague, Czech Republic
Dmitriy Voronin
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Lucia Votano
Laboratori Nazionali di Frascati dell’INFN, Roma, Italy
Caishen Wang
Dongguan University of Technology, Dongguan, China
Chung-Hsiang Wang
National United University, Miao-Li
En Wang
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Hanwen Wang
Institute of High Energy Physics, Beijing, China
Jiabin Wang
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Jun Wang
Sun Yat-Sen University, Guangzhou, China
Li Wang
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Institute of High Energy Physics, Beijing, China
Meng Wang
University of South China, Hengyang, China
Meng Wang
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Mingyuan Wang
Institute of High Energy Physics, Beijing, China
Qianchuan Wang
Wuhan University, Wuhan, China
Ruiguang Wang
Institute of High Energy Physics, Beijing, China
Sibo Wang
Institute of High Energy Physics, Beijing, China
Tianhong Wang
Harbin Institute of Technology, Harbin, China
Wei Wang
Sun Yat-Sen University, Guangzhou, China
Wenshuai Wang
Institute of High Energy Physics, Beijing, China
Xi Wang
College of Electronic Science and Engineering, National University of Defense Technology, Changsha, China
Yangfu Wang
Institute of High Energy Physics, Beijing, China
Yaoguang Wang
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Yi Wang
Institute of High Energy Physics, Beijing, China
Yi Wang
Tsinghua University, Beijing, China
Yifang Wang
Institute of High Energy Physics, Beijing, China
Yuanqing Wang
Tsinghua University, Beijing, China
Yuyi Wang
Tsinghua University, Beijing, China
Zhe Wang
Tsinghua University, Beijing, China
Zheng Wang
Institute of High Energy Physics, Beijing, China
Zhimin Wang
Institute of High Energy Physics, Beijing, China
Apimook Watcharangkool
National Astronomical Research Institute of Thailand, Chiang Mai, Thailand
Wei Wei
Institute of High Energy Physics, Beijing, China
Wei Wei
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Yadong Wei
Dongguan University of Technology, Dongguan, China
Yuehuan Wei
Sun Yat-Sen University, Guangzhou, China
Zhengbao Wei
Guangxi University, Nanning, China
Liangjian Wen
Institute of High Energy Physics, Beijing, China
Jun Weng
Tsinghua University, Beijing, China
Christopher Wiebusch
III. Physikalisches Institut B, RWTH Aachen University, Aachen, Germany
Rosmarie Wirth
Institute of Experimental Physics, University of Hamburg, Hamburg, Germany
Bi Wu
Sun Yat-Sen University, Guangzhou, China
Chengxin Wu
Sun Yat-Sen University, Guangzhou, China
Diru Wu
Institute of High Energy Physics, Beijing, China
Qun Wu
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Yinhui Wu
Institute of High Energy Physics, Beijing, China
Yiyang Wu
Tsinghua University, Beijing, China
Zhaoxiang Wu
Institute of High Energy Physics, Beijing, China
Zhi Wu
Institute of High Energy Physics, Beijing, China
Michael Wurm
Institute of Physics and EC PRISMA+, Johannes Gutenberg Universität Mainz, Mainz, Germany
Jacques Wurtz
IPHC, Université de Strasbourg, CNRS/IN2P3, F-67037 Strasbourg, France
Dongmei Xia
Chongqing University, Chongqing, China
Shishen Xian
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, China
Ziqian Xiang
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Fei Xiao
Institute of High Energy Physics, Beijing, China
Pengfei Xiao
Institute of High Energy Physics, Beijing, China
Xiang Xiao
Sun Yat-Sen University, Guangzhou, China
Wei-Jun Xie
Institute of Physics, National Yang Ming Chiao Tung University, Hsinchu
Xiaochuan Xie
Guangxi University, Nanning, China
Yijun Xie
Institute of High Energy Physics, Beijing, China
Yuguang Xie
Institute of High Energy Physics, Beijing, China
Zhao Xin
Institute of High Energy Physics, Beijing, China
Zhizhong Xing
Institute of High Energy Physics, Beijing, China
Benda Xu
Tsinghua University, Beijing, China
Cheng Xu
University of South China, Hengyang, China
Donglian Xu
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Fanrong Xu
Jinan University, Guangzhou, China
Jiayang Xu
Institute of High Energy Physics, Beijing, China
Jinghuan Xu
Guangxi University, Nanning, China
Meihang Xu
Institute of High Energy Physics, Beijing, China
Shiwen Xu
Institute of High Energy Physics, Beijing, China
Xunjie Xu
Institute of High Energy Physics, Beijing, China
Yin Xu
Nankai University, Tianjin, China
Yu Xu
Sun Yat-Sen University, Guangzhou, China
Jingqin Xue
Institute of High Energy Physics, Beijing, China
Baojun Yan
Institute of High Energy Physics, Beijing, China
Qiyu Yan
University of Chinese Academy of Sciences, Beijing, China
University of Warwick, University of Warwick, Coventry, CV4 7AL, United Kingdom
Taylor Yan
Suranaree University of Technology, Nakhon Ratchasima, Thailand
Xiongbo Yan
Institute of High Energy Physics, Beijing, China
Yupeng Yan
Suranaree University of Technology, Nakhon Ratchasima, Thailand
Changgen Yang
Institute of High Energy Physics, Beijing, China
Chengfeng Yang
Sun Yat-Sen University, Guangzhou, China
Fengfan Yang
Institute of High Energy Physics, Beijing, China
Jie Yang
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Lei Yang
Dongguan University of Technology, Dongguan, China
Pengfei Yang
Sun Yat-Sen University, Guangzhou, China
Xiaoyu Yang
Institute of High Energy Physics, Beijing, China
Yifan Yang
Université Libre de Bruxelles, Brussels, Belgium
Yixiang Yang
Institute of High Energy Physics, Beijing, China
Zekun Yang
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Haifeng Yao
Institute of High Energy Physics, Beijing, China
Jiaxuan Ye
Institute of High Energy Physics, Beijing, China
Mei Ye
Institute of High Energy Physics, Beijing, China
Ziping Ye
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, China
Frédéric Yermia
Nantes Université, IMT Atlantique, CNRS/IN2P3, Nantes, France
Jilong Yin
Institute of High Energy Physics, Beijing, China
Weiqing Yin
Institute of High Energy Physics, Beijing, China
Xiaohao Yin
Sun Yat-Sen University, Guangzhou, China
Zhengyun You
Sun Yat-Sen University, Guangzhou, China
Boxiang Yu
Institute of High Energy Physics, Beijing, China
Chiye Yu
Dongguan University of Technology, Dongguan, China
Chunxu Yu
Nankai University, Tianjin, China
Hongzhao Yu
Institute of High Energy Physics, Beijing, China
Peidong Yu
Institute of High Energy Physics, Beijing, China
Xianghui Yu
Nankai University, Tianjin, China
Zeyuan Yu
Institute of High Energy Physics, Beijing, China
Zezhong Yu
Institute of High Energy Physics, Beijing, China
Cenxi Yuan
Sun Yat-Sen University, Guangzhou, China
Chengzhuo Yuan
Institute of High Energy Physics, Beijing, China
Zhaoyang Yuan
Institute of High Energy Physics, Beijing, China
Zhenxiong Yuan
Tsinghua University, Beijing, China
Noman Zafar
Pakistan Institute of Nuclear Science and Technology, Islamabad, Pakistan
Kirill Zamogilnyi
Lomonosov Moscow State University, Moscow, Russia
Jilberto Zamora
Universidad Andres Bello, Fernandez Concha 700, Chile
Vitalii Zavadskyi
Joint Institute for Nuclear Research, Dubna, Russia
Fanrui Zeng
Shandong University, Jinan, China, and Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao, China
Shan Zeng
Institute of High Energy Physics, Beijing, China
Tingxuan Zeng
Institute of High Energy Physics, Beijing, China
Liang Zhan
Institute of High Energy Physics, Beijing, China
Yonghua Zhan
Sun Yat-Sen University, Guangzhou, China
Aiqiang Zhang
Tsinghua University, Beijing, China
Bin Zhang
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Binting Zhang
Institute of High Energy Physics, Beijing, China
Feiyang Zhang
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Han Zhang
Institute of High Energy Physics, Beijing, China
Haosen Zhang
Institute of High Energy Physics, Beijing, China
Honghao Zhang
Sun Yat-Sen University, Guangzhou, China
Jialiang Zhang
Nanjing University, Nanjing, China
Jiawen Zhang
Institute of High Energy Physics, Beijing, China
Jie Zhang
Institute of High Energy Physics, Beijing, China
Jingbo Zhang
Harbin Institute of Technology, Harbin, China
Junwei Zhang
Guangxi University, Nanning, China
Lei Zhang
Nanjing University, Nanjing, China
Ping Zhang
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Qingmin Zhang
Xi’an Jiaotong University, Xi’an, China
Rongping Zhang
Institute of High Energy Physics, Beijing, China
Shiqi Zhang
Sun Yat-Sen University, Guangzhou, China
Shuihan Zhang
Institute of High Energy Physics, Beijing, China
Siyuan Zhang
Guangxi University, Nanning, China
Tao Zhang
School of Physics and Astronomy, Shanghai Jiao Tong University, Shanghai, China
Xiaomei Zhang
Institute of High Energy Physics, Beijing, China
Xin Zhang
Institute of High Energy Physics, Beijing, China
Xu Zhang
Institute of High Energy Physics, Beijing, China
Xuantong Zhang
Institute of High Energy Physics, Beijing, China
Yibing Zhang
Institute of High Energy Physics, Beijing, China
Yinhong Zhang
Institute of High Energy Physics, Beijing, China
Yiyu Zhang
Institute of High Energy Physics, Beijing, China
Yongpeng Zhang
Institute of High Energy Physics, Beijing, China
Yu Zhang
Institute of High Energy Physics, Beijing, China
Yuanyuan Zhang
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, China
Yumei Zhang
Sun Yat-Sen University, Guangzhou, China
Zhenyu Zhang
Wuhan University, Wuhan, China
Zhijian Zhang
Dongguan University of Technology, Dongguan, China
Jie Zhao
Institute of High Energy Physics, Beijing, China
Runze Zhao
Institute of High Energy Physics, Beijing, China
Shujun Zhao
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, China
Tianhao Zhao
Institute of High Energy Physics, Beijing, China
Hua Zheng
Dongguan University of Technology, Dongguan, China
Yangheng Zheng
University of Chinese Academy of Sciences, Beijing, China
Li Zhou
Institute of High Energy Physics, Beijing, China
Shun Zhou
Institute of High Energy Physics, Beijing, China
Tong Zhou
Institute of High Energy Physics, Beijing, China
Xiang Zhou
Wuhan University, Wuhan, China
Xing Zhou
Institute of High Energy Physics, Beijing, China
Jingsen Zhu
Sun Yat-Sen University, Guangzhou, China
Kangfu Zhu
Xi’an Jiaotong University, Xi’an, China
Kejun Zhu
Institute of High Energy Physics, Beijing, China
Zhihang Zhu
Institute of High Energy Physics, Beijing, China
Bo Zhuang
Institute of High Energy Physics, Beijing, China
Honglin Zhuang
Institute of High Energy Physics, Beijing, China
Liang Zong
Tsinghua University, Beijing, China
Jiaheng Zou
Institute of High Energy Physics, Beijing, China
Abstract
Over 25,600 3-inch photomultiplier tubes (PMTs) have been instrumented for the central detector of the Jiangmen Underground Neutrino Observatory. Each PMT is equipped with a high-voltage divider and a frontend cable with waterproof sealing. Groups of sixteen PMTs are connected to the underwater frontend readout electronics via specialized multi-channel waterproof connectors. This paper outlines the design and mass production processes for the high-voltage divider, the cable and connector, as well as the waterproof potting of the PMT bases. The results of the acceptance tests of all the integrated PMTs are also presented.
1 Introduction
The Jiangmen Underground Neutrino Observatory (JUNO) is a particle physics experiment currently under construction, aiming for neutrino studies and exotic searches [1, 2, 3]. The central detector [4] is made of a 20 kiloton liquid scintillator target surrounded by 17,612 20-inch [5] and 25,600 3-inch photomultiplier tubes (PMTs) [6], all of which are submerged in water at a depth of up to 43.5 m along with their frontend electronics.
By virtue of their operation in single photon-counting mode at reactor antineutrino energies, the 3-inch PMT system provides a linear reference for calibrating any instrumental non-linearities in the charge response of the 20-inch PMTs [7, 8]. The 3-inch PMTs also improve the energy resolution [9], enhance event reconstruction [10], aid cosmic muon tracking [11], and support several physics measurements, including proton decay searches [12], and the measurement of the solar oscillation parameters and , for which they provide a semi-independent determination [13].
The system is divided into 200 readout units of 128 PMTs each. A diagram of one such unit is shown in Fig. 1. Each PMT is equipped with a high voltage (HV) divider and a coaxial cable that can be 5 m or 10 m in length and that transmits both the HV and the signal. The PMT, HV divider, and one end of the cable are soldered and waterproofed. The other ends of sixteen cables are connected to a multichannel connector, each consisting of a plug and receptacle pair. These connectors are linked to an underwater stainless steel box (UWB), housing the 128-channel frontend power supply and readout electronics. This paper provides details on the instrumentation of the 3-inch PMTs, encompassing the design and production of the HV divider, cable, and connector, as well as the waterproof potting and the acceptance tests. The production and characterization of 26,000 (including 400 spares) 3-inch PMTs was reported in Ref. [6]. The electronics design was outlined in Ref. [3, 14, 15].
The rest of this paper is structured as follows. In Sec. 2 and 3, the research and development, as well as the manufacturing process of the HV divider, cable, and connector, are presented. The waterproof potting is detailed in Sec. 4. Section 5 covers the acceptance tests following potting. Section 6 provides a summary.
Unless otherwise specified, all references to PMTs in the subsequent sections pertain to the 3-inch PMTs.
Figure 1: Diagram of a readout unit, consisting of 128 3-inch PMTs linked to underwater electronics through eight 16-channel connectors.
2 HV Divider
The HV divider follows a positive HV scheme. The HV ratio for the 11 series of dynodes is optimized for collection efficiency and timing measurements [16, 17, 18]. The resistor ratio for the first three dynodes is 3:2:1, and the other resistors are equal in value to the third one, 15 M (Fig. 2 (a)). The total resistance of 210 M and a working HV generally less than 1,300 V result in a gain of approximately and a direct current (DC) less than 6.5 µA. In consideration of safety and reliability requirements, the HV divider components include wire resistors, as depicted in Fig. 2 (b), due to superior voltage and thermal tolerance. The main challenge lies in the dense packing of components on the small PCB, which is designed to be soldered directly to the PMT pins and is constrained to have a similar size as the 48 mm diameter of the PMT glass bulb neck. Considering the re-evaluated requirements of the 3-inch PMT system for JUNO, the reliability requirement for the HV divider was slightly relaxed compared to the 20-inch PMT HV divider [5], allowing for the use of surface-mounted components. The final design is illustrated in Fig. 2 (c), where resistors of type 1206 with a maximum of 200 V and a 0.25 W power rating are selected, along with ceramic DC capacitors. The failure rate of the HV divider, as determined from reliability of individual components, is less than 0.05% per year.
The HV divider is designed to achieve a maximum working HV of 2,000 V using a single layer of standard FR-4 [19] PCB material with a thickness of 1.2 mm and a diameter of 47 mm, in compliance with the radioactivity requirements of JUNO [20]. The PCB design ensures a clearance exceeding 1.6 kV/mm. Testing was performed on a few random samples of the final products, where HV dividers were subjected to several weeks of operation up to 2,400 V with satisfactory performance, and a maximum of 3,000 V was applied for a short period without failure. The flasher rate of the HV divider was verified to be less than 1 mHz, as detailed in Ref. [21]. To mitigate potential risks of flasher occurrences during production and long-term operation, black polyurethane was used to cover the divider PCB after it has been soldered to the PMT pins, as will be discussed in Sec. 4. A temperature cycling test was successfully conducted on 20 samples for 10 cycles without any failures, subjected to temperatures ranging from 5∘C to 55∘C for three hours per cycle. The signal quality coupled with the 3-inch PMT was further assessed in Ref. [22, 23], with additional results to be discussed in the following sections.
Viking [24] surface bounding resistors of type 1206 and Murata [25] ceramic capacitors were chosen for the design. In adherence to the JUNO divider PCB procedure [21], mass production was executed after a successful small batch production of 100 pieces, all of which passed testing. All PCBs were soldered at the factory following cleaning at the designated soldering points. Subsequently, a visual inspection of the soldering points was performed. After another round of cleaning, each PCB underwent a 30-second maximum HV test at 2,000V before being soldered to the PMT and the cable. In total, 26,500 divider boards were produced which includes 900 spares. Acceptance tests were conducted during integrating with PMTs, as discussed in Sec. 4.2.
Figure 2: (a) Schematic design of the 3-inch PMT HV divider. R1 to R11 are divider resistors, each with a resistance of 15 M. K denotes the cathode, DY1 to DY10 represent the ten dynode stages, and P indicates the anode. C1 to C4 are decoupling capacitors, each with a capacitance of 10 nF, used to maintain the signal linearity. R12 to R14 are damping resistors, each with a resistance of 56 , to prevent signal ringing. R15 (100 k) and C5 (4.7 nF) together form a low-pass filter. R16 is a matched resistor with a resistance of 56 . (b) Initial version of the divider with wire resistors. (c) Final version of the divider with surface-mounted resistors.
3 Cable and Connector
3.1 Design
To reduce costs and facilitate the installation of the JUNO detector, a novel 16-channel waterproof connector and cable was designed in collaboration with AXON’ company [26]. The connector’s schematic design is shown in Fig. 3 and consists of two parts: a plug that is attached to 16 PMTs through frontend cables, and a receptacle that is mounted on the UWB. Waterproof sealing between the 16 cables is achieved through injection molding in the plug and the application of adhesive in the receptacle. A double-sealing design ensures tightness when the plug is connected to the receptacle.
The receptacle is approximately 25 cm in length and contains 16 specialized micro coaxial connector plug pins. These connector plugs are mounted on a special thermoplastic shell, which is subsequently inserted into the stainless steel shell of the receptacle. This design ensures that the cable grounding remains isolated from the receptacle shell, thus meeting JUNO’s requirement of signal grounding to be separated from grounding of the JUNO mechanical structure. The end of the receptacle shell is sealed with epoxy, providing the receptacle with independent waterproof capability, regardless of whether the plug is connected. Eight receptacles are mounted on the lid of the UWB. Even if one plug fails, the receptacle is designed to prevent water from entering the UWB, thereby protecting the internal electronics.
A custom-made RG178 coaxial cable with a 50 impedance was also developed. The outer layer of the cable is made of High-Density Polyethylene (HDPE) with a diameter of 2.1 mm, designed to prevent fluorescence and ensure long-term reliability in ultra-pure water environment. As an additional safeguard, a water-blocking powder, developed by AXON’ company, was incorporated into the grounding layer. If the cable jacket is compromised, water would penetrate this layer, activating this barrier and preventing further water ingress. A water pressure test demonstrated that after one hour under 5 bar pressure, water infiltration was restricted to only 5 cm.
During the commissioning phase of the JUNO detector, a few weeks after the start of water filling, a significant number of channels experienced HV breakdown. This issue was traced to air diffusing from the connectors toward the ultra-pure degassed water, resulting in a pressure drop within the connector assembly. To address this problem, a continuous flow of nitrogen gas was introduced into the water circulation system, effectively resolving the issue. The 3-inch PMT system was integrated into detector operation following the completion of liquid scintillator filling. The operating power of each PMT group is being continuously monitored and has remained stable. Further details and comprehensive analysis of this incident will be provided in an upcoming publication.
Figure 3: (a) Schematic design of the 16-channel connector consisting of a plug (left part) and a receptacle (right part). The red dots and lines indicate water-resistant O-rings or gaskets. Due to intellectual property rights, further details cannot be disclosed. (b) A picture of a connector set with cables.
3.2 Signal Performances
The quality of the PMT signal was examined using a prototype of the connector and cables, as detailed in Ref. [23]. Typical signals for the single photoelectron (SPE) with a gain of 3106 are displayed in Fig. 4 (a). Small reflections following the main signal, attributed to a slight impedance mismatch not to the connector but in the hand-made HV-signal splitter, were observed. This issue was resolved in the final version of the splitter as reported in Ref. [15]. To compensate for the loss of PMT gain due to cable attenuation, an additional 23 V or 45 V voltage was applied for 5 m or 10 m cables, respectively. The observed amplitude, rise time, fall time, and Full Width at Half Maximum are listed in Tab. 1.
Figure 4: (a) An average PMT signal for the single photoelectron using a 5 m (red) or 10 m (blue) cable. The reflection, resulting from a slight impedance mismatch in the prototype HV splitter, was addressed and resolved in the final version of the splitter [15]. (b) The average PMT signal with a large light intensity of approximately 4,000 photoelectrons.
Table 1: Measured features of PMT SPE signals with cable length 5 m and 10 m at a gain of . Errors are statistical only.
Gain
5 m
10 m
Amplitude (mV)
Rise time (ns)
Fall time (ns)
FWHM (ns)
The measured PMT charge is not proportional to the number of incident photoelectrons due to both internal factors, such as the space charge effect, and external factors, such as the divider current [27]. Fractional loss of charge, known as nonlinearity, was observed in the JUNO 3-inch PMTs and reported in Ref. [28]. The nonlinearity remains under 5% below 750 PEs but increases significantly above that due to the low divider current (less than 6.5 A). An average signal of approximately 4,000 incident PEs obtained with high-intensity LED lights is shown in Fig. 4 (b). The integrated charge corresponds to 2,342 PEs when comparing with the single PE signal in Fig. 4 (a), which represents about 40% charge nonlinearity. However, this nonlinearity is not an issue for JUNO, because the 3-inch PMTs operate in SPE regime for most physics cases.
The crosstalk for a complete assembly of cables, plugs and receptacles was assessed using an adapted network analyzer (Agilent E5062A) [29]. Considering two adjacent channels as the extreme case, in the frequency range of interest from 100 MHz to 350 MHz,
an S-parameter analysis verified that the analog charge crosstalk between different channels is less than 0.3%.
3.3 Long Term Validation
The JUNO 3-inch PMT system was designed to achieve a channel failure rate lower than 10% over a 6-year period, equally distributed among the PMTs and their readout electronics. This translates into an estimated Mean Time Between Failure (MTBF) exceeding 120 years, or equivalently, a Failure In Time (FIT) value below 1,000. This requirement applies directly to the cables and connectors that interface with the photomultipliers. Estimating such a low failure rate is challenging when dealing with a limited number of components over a short testing period. For instance, evaluating the reliability of two underwater connector plug and receptacle systems would necessitate a 40-year assessment.
In ultra-pure water, the primary cause of long-term failure is chemical degradation, which compromises the protective barriers against water ingress such as cable jackets and connector gaskets. The test duration can be accelerated by subjecting the system to increased temperatures in a controlled setting. Compared to JUNO’s operational water temperature of C, raising the temperature by an additional C can expedite the testing process by a factor of approximately 2.5. Thus, a test conducted at C for 12 months simulates the effect of 40 years of ageing at the normal operational temperature of C.
To conduct the test, two pressurized water tanks were developed and equipped - one located in France and the other in Chile. These tanks allowed for the adjustment of water temperature and pressure from 1 to 10 bars to simulate various depths in JUNO. In the test conducted in France, four underwater connectors with 10-meter long cables were mounted on their receptacles. An electric potential ranging from 0 to 2 kV was applied to each cable, with the leakage current monitored at the micro-ampere level, allowing for the detection of any minor water leaks on the cable jacket or at the connector level. In Chile’s test, five receptacles and three mechanical models were mounted on the UWB to evaluate long-term water leakage aging under conditions involving heat and pressure.
After the facilities were qualified, the following test campaigns were carried out:
A first period of 12 months between December 2020 and November 2021, with overpressure cycling between 5 to 10 bars and temperature varying between C and C in France;
A second period of 4 months between December 2021 and March 2022, with overpressure between 1 to 2 bars, mimicking shallow water, and a temperature varying between C and C in France.
Five receptacles with water temperature 88.7∘C and pressure from 4 to 9 bars, average 7.75 bars were tested in 2021 in Chile.
During the two testing phases conducted in France, spanning over 11,000 hours, no leaking current was observed, and both the cable and connector functioned perfectly, achieving the required reliability with a FIT value smaller than 500. During the destructive test conducted in Chile, five receptacles were aged to the equivalent of over 120 years and exhibited no leaks, yielding a FIT value of 350.
3.4 Production and Quality Inspection in Axon’
Approximately 200 km of cables were produced and thoroughly tested, with 100% of the insulation verified through a 3,000 V spark test. After the cabling and overmolding processes were finalized, each plug was tested with 1.8 kV of HV for 70 seconds to confirm functionality. Subsequently, the plugs were equipped with stainless steel accessories that matched those of the receptacle. The sealing was achieved using a double sealing design.
The sealed cables and plugs were placed into a stainless steel tank filled with ultrapure water, as shown in Fig. 5. Humidity test paper was also added. The tank was closed, and the water pressure was raised to 0.75 MPa, maintained for two hours. Once removed from the water tank and dried, the water sealing was verified through visual inspection of the humidity test paper and a subsequent HV test. All cables and connectors underwent testing. Those passing the tests were then used to create the required 10 m and 5 m segments.
Figure 5: Leak test in a high-pressure water tank at AXON’. The dual-end plugs were sealed with fake receptacles and directly put into water. The receptacles were mounted on the lid of the tank.
For each batch, the first and last receptacles were mounted on the lid of the water tank with the 16 pins to be submerged in water (Fig. 5). If these two receptacles passed the test, the remaining receptacles from the batch proceeded to the next step, which involved attaching the micro coaxial connector to the other end of the cable that connects to the electronics in the UWB. In the case of failure, a more thorough inspection and additional destructive pressure testing were conducted on the batch in question. Finally, the receptacle underwent the standard HV test at 1.8 kV for 70 seconds.
In total, 825 plugs of 10 meters in length and 825 plugs of 5 meters in length, along with 1,650 receptacles (including 50 spares), were produced and delivered to JUNO.
3.5 Acceptance Tests by JUNO
A warehouse designated for the JUNO 20-inch PMT mass testing [5] was also used for the storage and acceptance tests of the 3-inch PMT cables and connectors. The initial step involved visual inspections to detect any visible damage or defects, with particular attention paid to scratches on the sealing surfaces of the connectors. Electrical tests followed, assessing the resistance and electrical conduction of the core wire and the braid grounding wire, as well as verifying the channel sequence from 1 to 16. Additionally, HV testing (1.6 kV, 5 minutes) was carried out to determine the cables’ ability to withstand voltage stress and confirm their safe operation under normal conditions. Since only two receptacles from each batch underwent destructive tests underwater at Axon’, a supplementary leak test with SF6 gas was conducted for 10% of the receptacles. Each selected sample was sealed with a fake plug using an axial O-ring, evacuated to below 50 Pa, maintained for one minute, and then filled with SF6 gas at a pressure about 25% higher than standard atmospheric pressure. An SF6 detector was used outside the receptacle to detect any leaked gas with a sensitivity of 10Pam3/s [30]. No leakage was detected in any of the tested receptacles.
Figure 6: An SF6-based leakage test facility during inspection. Two receptacles undergo evacuation and are subsequently filled with SF6 gas to perform the leakage test.
4 Integration and Waterproof Potting
4.1 PMT Grouping
The 26,000 qualified and delivered bare PMTs [6] were assembled into groups of 16 according to their working HV at a gain of . Additionally, to minimize the risk of PMT implosion underwater, these PMTs were grouped according to their weights and installed at similar depths in the JUNO detector.
Grouping by HV
The operating HV required to achieve a gain of for each PMT was determined by the manufacturer, with the distribution shown in Fig. 7. The voltage ranges from 900 V to 1,300 V. PMTs with working voltage outside of this range were excluded from mass production [6]. The voltages were categorized into 16 groups, each spanning 25 V, with a maximum gain difference of 15% to 20% within each group.
Figure 7: Distribution of working voltage at gain of PMTs.
Grouping by weight
PMTs are made of quartz glass with an average thickness between 1 mm and 2 mm. During long-term operation in water at a 42 m depth, buckling failure may happen due to large external pressure. For an ideal spherical shell, the classical critical buckling pressure can be expressed as [31]
where is the Poisson’s ratio, is the tensile modulus, and
and are the radius and thickness of the glass shell, respectively. Since and are constants for the PMT glass shell, and the shape () is almost fixed, the critical pressure is sensitive to the thickness of the glass and thus to the weight.
A dedicated study was done to find the dependence of the critical pressure on the weight of the PMT. Ten PMTs were randomly selected from an early production batch and their weights were measured. All pins on the bottom of the PMT were sealed with polyurethane and each PMT was individually put into a water tank. The water pressure was increased with a water pump at a speed of about 0.5 MPa/min. When the PMT glass imploded, the water pressure was recorded.
Critical pressures at different PMT weights are shown in the bottom panel of Fig. 8. A positive correlation can be seen. However, there is a large spread from 1.5 MPa to 2.5 MPa in the 110 g to 115 g range, which can be attributed to variations in the glass shell thickness, as implosions typically initiate at points with the smallest thickness or highest stress.
A parabolic function (red dashed curve) is used to fit the critical pressure distribution, but with large fluctuation due to the special shape (not ideal spherical) and non-uniform thickess of the glass shell, as well as the weight of electric components inside the shell.
Figure 8: The top panel shows the measured weight distribution for all 26,000 PMTs. W1, W2, and W3 represent three groups of PMTs according to their weights. The blue points in the bottom panel are the measured critical pressures underwater as a function of the PMT weight. The red dashed line is the best fit parabola.
The weights of all 26,000 PMTs were measured and the resulting distribution is shown in the top panel of Fig. 8.
They were categorized into three classes based on their weights and installed at maximum water depths of 10 m, 25 m, and 42 m, respectively, as detailed in Tab. 2. The critical pressure for each class is approximately five times greater than the expected water pressure. Only 3.8% of PMTs fall into class W1 and were installed near the top of the JUNO detector at less than 10 m depth. About 40.5% PMTs, belonging to class W2, were installed on the upper hemisphere of the JUNO detector at depths smaller than 25 m. The remaining PMTs, comprising more than half of the total, belong to class W3 and were installed mostly on the lower hemisphere.
Table 2: Classification of PMTs by weight.
Class
Fraction
Weight
Installation depth
Water pressure in the leakage test
W1
3.8%
[90, 105] g
10 m
0.20 MPa
W2
40.5%
[105, 116] g
25 m
0.40 MPa
W3
55.7%
[116, 150] g
42 m
0.65 MPa
In total, 48 groups were formed, comprising three weight-based groups and 16 HV-based groups. During the potting process, PMTs were selected from groups that were as homogeneous as possible. The installation depth was determined based on the lightest PMT in each assembly.
4.2 Waterproof PMT Potting
The schematic design of the waterproof PMT potting is illustrated in Fig. 9. The PMT pins, the HV divider, and the cable are soldered and enclosed in an ABS plastic shell. This shell is filled with polyurethane to protect the electrical components from water and to block any light that might be emitted spontaneously by electrical discharges. The seal between the shell and the PMT glass is achieved using butyl tape, which is encompassed by a shrinkable tube and finished with low-pressure injection molding at the end.
Figure 9: Schematic design of waterproof PMT potting.
As a final verification of the waterproof potting design, 50 samples were produced using PMT glass bulbs or discarded PMTs along with HV dividers and short cables. The cable ends were sealed with epoxy, and all samples were submerged in a water tank in four separate batches, subjected to a pressure of 0.6 MPa for a minimum of 7 days. A potential difference of 2,000 V was applied through the cable ends, and currents were measured both before and after the high-pressure water test. The observed differences ranged from -40% to 70%, which can be attributed to variations in PMT exposure to light. If water had penetrated into the PMT pins or the divider, the current would have increased significantly, by at least an order of magnitude.
The primary steps involved in the potting are illustrated in Fig. 10 (a). An injection-molded ABS plastic shell was designed to provide mechanical support. Initially, the cable was threaded through a bottom entrance with a diameter of 2.2 mm, ensuring 28 cm of the cable remained inside the shell. Following this, epoxy was injected from the top to offer a bonding length greater than 8 cm, serving as the inner seal. The epoxy used is specialized for HDPE, featuring a bonding strength of 1.6 MPa. The sealing was further reinforced outside the shell’s tail through low-pressure injection molding with a polyamide and polyolefin mixture, also customized for HDPE.
Figure 10: (a) Major steps of waterproof PMT potting: 1, cable sealing with epoxy; 2, cable sealing with injection molding; 3, soldering HV divider on the bare PMT; 4, assembly of butyl tape around the PMT; 5, filling polyurethane; 6, inverting the PMT; 7, covering the shrinkable tube. The PMT signals were verified after step 3 and step 7 with HV. (b) A group of 16 PMTs with cables and the connector after waterproof potting.
Subsequently, the PMT was mounted and soldered onto the HV divider along with the cable. A set of 16 PMTs with plugs were then transferred to a dark box, where each channel was connected to a splitter box to separate the HV and the signal. A nominal HV was applied to each channel, and the current was monitored using an ammeter. Utilizing low-intensity LED lights inside the dark box, PMT signals, predominantly SPEs, were captured by an oscilloscope. The amplitude was required to be within 50% of the average. If a significant deviation was detected, the PMT was replaced, and the process was repeated for a new PMT.
A fixed volume of a two-part polyurethane adhesive was mixed and filled into the potting shell. The adhesive did not completely fill the potting shell, leaving a few milliliters of air to allow for sealant expansion due to temperature changes. The PMT glass was encircled by a 5 cm wide layer of butyl tape near the bottom of the PMT and positioned in the shell, where a pedestal provided mechanical support. This butyl tape filled the gap between the PMT glass and the shell. The PMT with the potting shell was inverted to ensure complete coverage of all electrical components by the polyurethane, and it was left for 24 hours to allow the polyurethane to cure. A second layer of butyl tape was then added outside the shell as an outer seal, with a 2 cm overlap on the first layer. Finally, a plastic shrinkable tube was added to protect the layers of butyl tape. The fully potted PMTs were tested again to verify signal quality. A picture of an integrated group of 16 PMTs is shown in Fig. 10 (b).
4.3 Leakage Test
A large steel water tank was utilized to conduct a sampling test on potted PMTs to verify water leakage, as depicted in Fig. 11. A group of 16 PMTs was secured on a rack, with the plug sealed using two O-rings within a customized receptacle. Nine groups of PMTs of identical weight classification were simultaneously placed in the tank. The water pressure was elevated according to the weight specifications listed in Table 2 and maintained for 48 hours. Subsequently, the PMTs were removed from the tank, rinsed with pure water, and dried in an oven at 55oC. The PMTs were then subjected to HV testing, and signals were measured for the third time in the dark box.
Figure 11: (a) Water tank used for the PMT potting leak testing. (b) Nine groups of PMTs taken out of the tank after 48 hours. Each group of 16 PMTs was fixed on a plastic rack.
A total of 217 groups of PMTs, constituting approximately 13% of the entire collection, were randomly selected from all three weight classes and subjected to underwater testing.
No potting leaks were detected in any of the groups. Nevertheless, four PMT glasses fractured during the water pressure test. Of these, two belonged to weight class W1, and the other two to weight class W2. The failure rate was approximately 0.1%. These PMTs were subsequently replaced.
5 Acceptance Test of Potted PMTs
After potting, all PMTs underwent another round of testing to ensure their reliability and performance prior to being delivered to the JUNO site. This section provides details of the acceptance tests conducted.
5.1 Test Station
Figure 12: Schematics of a 128-channel PMT acceptance test station and an image of a 128-channel electronics test box.
To ensure efficiency and speed during testing, a test station consisting of two test benches was constructed at Guangxi University (GXU). Each test bench can simultaneously handle eight groups of PMTs. The acceptance test setup included a dark room where the PMT transportation and storage boxes can be placed. Without removing the PMTs from their boxes, they can be connected via feedthrough wall connectors to a 128-channel electronic test box. Figure 12 provides a schematic view of the complete acceptance test system and a photo of the electronics box. This test box is an early simplified prototype of the electronics system. It comprised a 128-channel front-end readout board called the ABC (ASIC Battery Card), connected to the photomultipliers through two 64-channel splitter boards [15]. The prototype ABC board was equipped with 8 CATIROC ASIC chips [14]. Each prototype splitter board was designed to receive four HV inputs from a commercial HV power supply while still handling 64 channels. The firmware and software were customized for the acceptance test scenarios.
5.2 Acceptance Criteria
The characterization of bare PMTs was conducted at the factory of Hainan Zhanchuang Photonics Technology Co., Ltd (HZC) [32], as noted in Ref. [6].
The primary objective of the acceptance tests was to verify the functionality of each PMT following waterproof potting. Consequently, three main performance parameters were established as acceptance criteria, based on experimental requirements detailed in Table 3. The consistency of the working HV is a crucial parameter because groups of potted PMTs share the same HV. When the average gain for each group of PMTs is , the gain of each PMT should should range between and . The SPE resolution of each PMT must be less than 45%, and the dark count rate (DCR) must be lower than 3 kHz.
Table 3: Acceptance test criteria.
Parameters
Requirements
Gain
SPE resolution
Dark rate
kHz
5.3 HV Calibration Testing
The working HV for the potted PMTs is higher than that of the bare tubes. This is because, firstly, the HV for potted PMT is supplied through an HV splitter. Each channel’s 20 M current-limiting resistor reduces the HV available to the tube by approximately 10%. Secondly, as discussed in Sec. 3.2, any attenuation in the long cable is compensated by applying a slightly higher HV. At the start of the acceptance test, the working HV for achieving a gain of (designated as GXU HV) was determined by systematically scanning various HV levels for small batches of PMTs.
This value was then compared with the working HV determined from bare tube testing by the PMT manufacturer HZC (referred to HZC HV, with each increment of 25 V represents a distinct HV level). This comparison resulted in a linear relationship, as depicted in Fig. 13.
Figure 13: Correlations between GXU and HZC HV measurements. The red line is the result of a simple linear fit to the data. (a) The cable length is 5 meters. (b) The cable length is 10 meters.
Utilizing this correlation, the working HV for each group of PMTs can be easily derived from their HZC HV values during acceptance testing. Subsequent tests confirmed the effectiveness of this procedure, as only a small number of PMT groups required further adjustments to their working HV during testing. These instances were often due to the challenge of assembling 16 PMTs with identical HV levels into a complete set during waterproof potting, resulting in PMTs with significantly different HV levels being included in the same group. Once all PMTs are installed in the JUNO detector, the scanning will be conducted again for every group to determine the precise working HV.
5.4 Test Results
As stated in Sec. 5.1, for the efficient batch acceptance testing of 26,000 PMTs, we employed two test benches capable of simultaneously measuring 128 PMTs per test setup. The PMTs were placed inside a storage box and shielded in a dark room with a black cloth. It was found that using PMT dark noise hits for gain and SPE resolution measurements was the optimal choice. Before collecting test data, the PMTs were placed in the dark room and powered at their working HV for more than eight hours to reduce and stabilize the DCRs. Forced trigger mode (pedestal mode) was used in the tests to establish the baseline level of each channel. Subsequently, self-triggered mode was utilized to capture the SPE signals generated by PMT dark count events, with a trigger threshold set at 870 digital to analog converter units (DACu), corresponding to about 0.3 PE.
Figure 14 and Fig. 15 show the pedestal and SPE distributions for a group of 16 PMTs, respectively. By fitting these distributions with Gaussian functions, the analog to digital converter units (ADCu) corresponding to their peak positions can be determined, referred to as and , respectively. Additionally, the sigma value of the Gaussian function is denoted as for the SPE distribution. Consequently, the gain of each PMT can be calculated using the following formula:
where is the calibration constant of each
CATIROC channel, which is ADCu/pC for a CATIROC preamplifier gain of 40. As a result, the estimated range for the ADC value () corresponding to a gain is approximately from 77 ADCu to 91 ADCu.
Figure 14: Pedestal charge spectra of a group of PMT.Figure 15: The SPE charge spectra of a group of PMTs.
In the analysis of the test data, to facilitate the comparison and evaluation of gain consistency within the same group of PMTs under a specified working HV, the average gain of the group was normalized to . The gains of PMTs in the same group were scaled by the same normalization factor. Subsequently, PMTs that did not meet the acceptance criteria (gain below or above ) were excluded and replaced with spare PMTs that aligned with the working HV.
The gain distribution of the measured PMTs as obtained from this analysis is shown in Fig. 16(a). The results presented in Fig. 16(b) demonstrate that the gain spread remains consistently around across different HV levels.
Figure 16: (a) Gain distribution of the measured PMTs. (b) Gain spread of PMTs across the 16 HV levels. The red line represents a constant fit to all points, illustrating that the spread remains consistently around 10% across all HV levels.
In addition, the SPE resolution was obtained as follows:
The SPE resolution results are presented in Fig. 17, indicating that the majority of PMTs exhibit a value below 45%, thereby satisfying the acceptance criteria. The SPE resolution was further compared with the results of the bare tube tests, and good consistency was observed.
Figure 17: (a) The SPE resolution of the measured PMTs. (b) The comparison of SPE resolution between the potted PMTs and the bare PMTs.Figure 18: The S-curve for one PMT. The number of triggers counted in 0.419 s as a funnction of the trigger threshold in DAC unit (a high DACu value corresponds to a low trigger threshold value in amplitude).
Figure 19: PMT dark rate distributions measured at the company (HZC) and during the acceptance tests at Guangxi University (GXU).
Another important acceptance parameter is the DCR, which is measured by the threshold scanning method. In this test, the number of trigger signals from each PMT channel in an electronics time period of 0.419 s is measured as a function of the trigger threshold in DAC units ranging from 0 to 870. The trigger values are in inverse order: a high DACu value corresponds to a low trigger threshold value in amplitude. The counts versus the discriminator threshold (also called S-curve) is shown in Fig. 18. The PMT DCR was obtained by calculating the trigger rate at threshold 870 DACu, which corresponds to 0.3 PE. Figure 19 shows the DCR distribution of the potted PMTs and the bare PMTs. The average DCR for the bare PMTs is approximately 510 Hz, with a significant tail in the distribution primarily attributed to insufficient time spent in the dark room under high voltage (HV) for the DCR to stabilize before testing. For instance, some PMTs were only subjected to a 2-hour HV load, and a DCR value below 3 kHz was considered acceptable. The potted PMTs benefited from sufficient cooling time with working HV of 8 hours, resulting in a significant reduction in the tail component and an average dark count rate of approximately 420 Hz, lower than that observed for the bare PMT test in HZC. However, upon comparing the main peak position of the dark count rate distribution, a significant shift is observed in the acceptance test, with the peak position approximately 100 Hz higher than that in the bare PMT test. Our investigation reveals that this additional contribution primarily comes from fluorescence emitted by the packaging foam employed to protect the PMTs. Subsequent tests conducted on small batches of PMTs demonstrated a difference of around 100 Hz in dark count rates between foam-protected and non-foam conditions. Given that the primary objective of the acceptance test is to validate the functionality of the potted PMTs, it was deemed unnecessary to conduct extensive testing without foam protection in order to ensure efficiency and prevent any damage during this process.
In total, 25,744 instrumented PMTs, corresponding to 1,609 groups, were tested. The rejected ratio of the acceptance test is about 0.7%. Table 4 summarizes the number of unqualified PMTs and the issues identified during testing. PMTs that did not pass the tests were removed and replaced with spare PMTs. The replaced PMTs were subsequently tested at the JUNO site, where they all met the acceptance criteria.
Table 4: The result of acceptance test.
Issue
Unqualified num.
Gain + Resolution + Dark count rate
1
Gain + Resolution
14
Gain + Dark count rate
0
Resolution + Dark count rate
0
Gain
58
SPE resolution
14
Dark rate
81
Short-circuit
9
Total
177
6 Summary
The instrumentation of the 3-inch PMTs of JUNO underwent rigorous testing at every stage of development and mass production. Each PMT was equipped with a divider capable of supplying up to 1,300 V across the various dynodes. Capacitors with wires and surface-mounted resistors were meticulously chosen to meet space constraints while ensuring reliability. A standard RG178 coaxial cable, available in 5 m or 10 m lengths, was used to transmit both HV and the signal. The cable jacket was customized to enhance waterproofing. A 16-channel waterproof connector, developed in collaboration with the Axon’ cable company, served as the interface between the PMTs and the electronics. Signal attenuation in the cable was adequately compensated by the voltage, and the observed analog crosstalk was below 0.3%. Two long-term underwater validations were conducted, demonstrating promising reliability with FIT values of less than 500 and 350, respectively, in line with JUNO’s requirements.
To integrate PMTs with HV dividers, cables, and connectors, the PMTs were grouped into more than 1,600 sets based on their weights and operating HV. Each group was connected to a single connector and arranged in the water pool from shallow to deep, according to weight, from lightest to heaviest. Waterproof potting of the PMTs was achieved through multiple redundant layers of sealants with mechanical support. A random selection of 217 groups was tested underwater, and no leakage was found in the PMT sealing.
All potted PMTs were tested with an early version of the electronics system. The SPE spectrum of each group was verified under the nominal threshold. A 10% standard deviation from the nominal 3106 gain was found. After potting, the SPE resolution remained unchanged, and there was no significant variation in the DCR. Only 0.7% of potted PMTs were replaced due to high DCRs, significant deviations from the target gain, poor charge resolution, or electrical short circuits. Altogether, 25,600 instrumented PMTs, along with an additional 144 spares, were delivered to JUNO.
7 Acknowledgement
We are grateful for the ongoing cooperation from the China General Nuclear Power Group.
This work was supported by
the Chinese Academy of Sciences,
the National Key R&D Program of China,
the Guangdong provincial government,
and the Tsung-Dao Lee Institute of Shanghai Jiao Tong University in China,
the Institut National de Physique Nucléaire et de Physique de Particules (IN2P3) in France,
the Istituto Nazionale di Fisica Nucleare (INFN) in Italy,
the Fond de la Recherche Scientifique (F.R.S-FNRS) and the Institut Interuniversitaire des Sciences Nucléaires (IISN) in Belgium,
the Conselho Nacional de Desenvolvimento Científico e Tecnològico in Brazil,
the Agencia Nacional de Investigacion y Desarrollo and ANID - Millennium Science Initiative Program - ICN2019_044 in Chile,
the European Structural and Investment Funds, the Czech Ministry of Education, Youth and Sports and the Charles University Research Centerin Czech Republic,
Deutsche Forschungsgemeinschaft (DFG), the Helmholtz Association, and the Cluster of Excellence PRISMA+ in Germany,
the Joint Institute of Nuclear Research (JINR) and Lomonosov Moscow State University in Russia,
MOST and MOE in Taipei,
the Program Management Unit for Human Resources & Institutional Development, Research and Innovation, Chulalongkorn University, and Suranaree University of Technology in Thailand,
the Science and Technology Facilities Council (STFC) in the UK,
University of California at Irvine and the National Science Foundation in the US.